Evaluate the limit as x turns to 1-

Lim [sin(x2-1) / e(Inx) - 1]



1
Expert's answer
2022-02-01T12:41:01-0500

"\frac{sin(x^2-1)}{e^{lnx}-1}=\frac{sin(x^2-1)}{x-1}"


using L'hopital's rule:

"(sin(x^2-1))'=2xcos(x^2-1)"

"(x-1)'=1"


"\displaystyle \lim_{x\to 1^-}\frac{sin(x^2-1)}{x-1}=\displaystyle \lim_{x\to 1^-}2xcos(x^2-1)=2"